For the follwoing date find:

30 20 35 17 50 54
64 29 34 23 20 20
21 50 26 26 20 48
20 30 23 29 17 17
20 18 23 36 20 20

1. The Mean=
2. The Median=
3. The Mode=
4. The Range =
5. The Standard Deviation is =
6. The Variance is=
7. Q1Q1 =
8. Q3Q3 =
9. IQR=
10. Upper Fence=
11. Lower Fence is =
12.If you graphed the data above, what shape would you expect the data to be?

  • Skewed-Right
  • Bell-Shaped
  • Skewed-Left
Answer #1

Solution:

For the given data, we have to find measures of central tendency and variation. The calculation table required for the calculation purpose is given as below:

No.

X

(X - mean)^2

1

17

136.1111889

2

17

136.1111889

3

17

136.1111889

4

18

113.7778489

5

20

75.11116889

6

20

75.11116889

7

20

75.11116889

8

20

75.11116889

9

20

75.11116889

10

20

75.11116889

11

20

75.11116889

12

20

75.11116889

13

21

58.77782889

14

23

32.11114889

15

23

32.11114889

16

23

32.11114889

17

26

7.111128889

18

26

7.111128889

19

29

0.111108889

20

29

0.111108889

21

30

1.777768889

22

30

1.777768889

23

34

28.44440889

24

35

40.11106889

25

36

53.77772889

26

48

373.7776489

27

50

455.1109689

28

50

455.1109689

29

54

641.7776089

30

64

1248.444209

Total

860

4592.666667

Mean

28.66667

We have

Sample size = n = 30

Mean = ∑X/n = 860/30 = 28.66667

Median = Middle most observation when data is in an increasing or decreasing order

Here, n = 30, so median is the mean of 15th and 16th observations.

Median = (23 + 23)/2 = 23

Median = 23

Mode = Observation with highest frequency

Mode = 20

Minimum = 17

Maximum = 64

Range = Maximum – Minimum = 64 – 17 = 47

Standard deviation = sqrt[∑(X – mean)^2/(n – 1)]

Standard deviation = sqrt[4592.666667/(30 – 1)]

Standard deviation = sqrt(158.3678161)

Standard deviation = 12.58442752

Variance = ∑(X – mean)^2/(n – 1)

Variance = 4592.666667/(30 – 1)

Variance = 158.3678161

Q1 = 20

Q3 = 34

(by using excel)

IQR = Q3 – Q1 = 34 – 20 = 14

Upper Fence = Q3 + 1.5*IQR = 34 + 1.5*14 = 55

Lower Fence = Q1 – 1.5*IQR = 20 - 1.5*14 = -1

Boxplot for the given data is given as below:

From above box plot, it is observed that the given data is skewed at right.